OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (phi):

equation of a circle is \[ (x-x_0)^2 + (y-y_0)^2 = r^2 \] where \( (x_0,y_0) \) is the center of the circle, and r is the radius

OpenStudy (anonymous):

i think it is (x – 2)2 + (y + 1)2 = 8

OpenStudy (phi):

close, but you want the radius squared on the right hand side. 8*8=64